# Causal Concept Embedding Models: Beyond Causal Opacity in Deep Learning

### v1

@@ 5.1 Key findings
We can also quantify the strength of the causal dependency between two nodes by computing the probability of necessity and sufficiency (PNS) [43]. In the figure, we represent the PNS w.r.t. the leaf node by colouring each node with a different shade of orange. This shows how, for Causal CEM, the attribute “heavy makeup” has the strongest impact on the leaf node. This high degree of causal transparency allows users to interpret Causal CEM’s inference and can be eventually exploited to identify potential biases, thereby supporting the assessment of the model’s counterfactual fairness. As a result, users can intervene directly on the causal structure of the decision-making process and remove biases using do-interventions, as shown in the next paragraph.
@@
checkmark dsprites celeba
Black box N/A N/A N/A
CBM $97.99{\scriptstyle\pm 5.64}$ $100.00{\scriptstyle\pm 0.70}$ $97.84{\scriptstyle\pm 2.13}$
CEM $102.58{\scriptstyle\pm 12.95}$ $100.00{\scriptstyle\pm 4.62}$ $106.00{\scriptstyle\pm 0.50}$
Causal CEM+CD $0.00{\scriptstyle\pm 0.00}$ $0.00{\scriptstyle\pm 0.00}$ $0.00{\scriptstyle\pm 0.00}$
Causal CEM $0.00{\scriptstyle\pm 0.00}$ $0.00{\scriptstyle\pm 0.00}$ $0.00{\scriptstyle\pm 0.00}$
Causal CEMs can make two causally-related variables causally independent by blocking all paths between these variables (Table 2)
The causal transparency of the proposed approach allows users to modify the model’s decision-making process (e.g., to de-bias the model’s inference) by using do-interventions. In particular, we can make two causally-related variables $i$ and $j$ causally independent by blocking all paths between the cause $i$ and the effect $j$. We experimentally verify this property and its impact by computing the Residual Concept Causal Effect i.e., the ratio between the Concept Causal Effect (CaCE) [44] obtained after and before blocking. The optimal value of this metric is zero, corresponding to perfect causal independence between $i$ and $j$ (i.e., the optimal value for a de-biasing operation). The results show that, in Causal CEMs, blocking a variable in the causal graph always yields a perfect Residual Concept Causal Effect of zero across all datasets. In contrast, applying the same procedure in CBMs leads only to a negligible reduction in the average causal effect to $3$ percentage points. CEMs not only fail to reduce the causal effect to zero but, in some cases, even increase the causal influence. These results underscore how Causal CEMs transparency enable users to manipulate the model’s decision-making process to achieve desired outcomes, as opposed to existing CBMs.
@@ 6 Discussion
Causal CEMs present substantial advantages compared to the state of the art. Compared to most causal feature-attribution methods (e.g., [13]), Causal CEMs focus on high-level human interpretable concepts. Causal CEMs differ from existing CBMs in their approach to intervention and causal relationships. The causal structure of the inferences that Causal CEMs deliver does not subsume weak forms of causal dependencies as in existing CBMs, where concepts are all direct causes of the target and causally independent of each other. This way, Causal CEMs enable human-in-the-loop corrections to mispredicted intermediate reasoning steps, boosting not just downstream accuracy after corrections but also the accuracy of the explanation provided for a specific instance. Closest to Causal CEMs in concept usage are post-hoc causal concept-based explainability techniques, like DiConStruct [22] and conceptual counterfactual explanations [25]. These methods build surrogate causal models to emulate a target black box model’s predictive behavior. However, as Rudin [45] notes, convergence in predictive behavior does not ensure structural similarity in decision-making processes, making surrogate models as explanatory proxies questionable. Causally interpretable by-design architectures, such as Causal CEMs, do not suffer from this issue.
@@
Our method’s limitations are mainly derived from limitations inherent to CBMs and causal reasoning. The quality of learnt causal graphs mainly depends on the quality of the dataset and its annotations. Missing and noisy labels might lead to suboptimal graphs. Similarly to generalised PGMs, Causal CEMs can be easily unfolded when the final graph is acyclic. Variables in cycles can still be inferred, but require special unfolding techniques which might be explored in future works. Finally, a Causal CEM’s learnt graph does not necessarily represent the causal mechanisms of the data-generating process, but rather that of the model’s inference. This makes our models suitable to be verified and controlled but not necessarily to understand their training data’s distribution.
Conclusion and impact
@@
Causal opacity represents a key open challenge at the intersection of deep learning, interpretability, and causality. Causal CEMs address this challenge by employing an architecture which makes the decision-making process causally transparent by design. This makes Causal CEMs reliable and verifiable compared to both usual DL architectures and standard (non-causal) CBMs. The results of our experiments show that Causal CEMs support the analysis of interventional and counterfactual scenarios—thereby improving the model’s causal interpretability and supporting the effective verification of its reliability and fairness—and enable human-in-the-loop corrections to mispredicted intermediate reasoning steps, boosting not just downstream accuracy after corrections, but also accuracy of the explanation provided for a specific instance. As a result, advancing this research line could significantly improve the reliability and verifiability of concept-based deep learning models, thus supporting their deployment in real-world applications.
Appendix A Architecture
Initialization of adjacency matrix

### v4

@@ 4.1 Key findings
In Causal CGM the decision-making process is causally interpretable and can be analysed by visualising the learnt causal graph and structural equations as in a structural causal model, as shown in Figure 6(a) for CelebA and in Figure  6(b) for Cifar10 (more results in App. H.2). The first image shows how Causal CGM exploited known biases in CelebA to infer facial attributes. For instance, Causal CGM predicts the attribute “wearing lipstick” directly from the attribute “attractive” and indirectly from attributes such as “smiling” or “high cheek”, all attributes that are known to be strongly correlated with each other in CelebA (Ramaswamy et al., 2021; Wang & Russakovsky, 2023). Similarly, in the CIFAR-10 dataset, where concepts are automatically extracted, the model learns meaningful relationships between concepts (Figure 6(b)), such as the presence of a “port” implying the presence of a “dock”, or the connection between a “beak” and a “bird”. We can also quantify the strength of the causal dependency between two nodes by computing the probability of necessity and sufficiency (PNS) (Pearl, 2022). In the figure, we represent the PNS w.r.t. the leaf node by colouring each node with a different shade of orange (more PNS of the other datasets in App. I). This shows how, for Causal CGM, the attribute “heavy makeup” has the strongest impact on the leaf node. This high degree of causal transparency allows users to interpret Causal CGM’s inference and can be eventually exploited to identify potential biases, thereby supporting the assessment of the model’s counterfactual fairness (an example of generated counterfactuals in Table  2). As a result, users can intervene directly on the causal structure of the decision-making process and remove biases using do-interventions, as shown in the next paragraph.
(a) CelebA
(b) Cifar10
Figure 6: Portion of the learnt causal graph and structural equations in CelebA (on the left) and Cifar10 (on the right). A node’s colour in the causal graph is proportional to the probability of necessity and sufficiency w.r.t. to the child nodes.
Table 2: Examples of counterfactuals generated on CelebA obtained via do-interventions (intervened variables are marked in red, changed variables are underlined).
Endogenous variables’ activations
Initial Predicted State Smiling, Not Attractive, Mouth Slig, High Cheek, Not Wearing Li, Not Heavy Make,
Not Wavy Hair, Not Big Lips, Not Oval Face, Not Makeup, Not Fem Model
What if I would wear lipstick? Smiling, Attractive, Mouth Slig, High Cheek, Wearing Li, Heavy Make,
Not Wavy Hair, Not Big Lips, Not Oval Face, Not Makeup, Not Fem Model
What if I would wear lipstick and also makeup? Smiling, Attractive, Mouth Slig, High Cheek, Wearing Li, Heavy Make,
Not Wavy Hair, Big Lips, Not Oval Face, Makeup, Fem Model
Causal CGMs can make two causally-related variables causally independent by blocking all paths between these variables (Table 3)
The causal transparency of the proposed approach allows users to modify the model’s decision-making process (e.g., to de-bias the model’s inference) by using do-interventions. In particular, it enables making two causally-related variables $i$ and $j$ independent by blocking all paths between the cause $i$ and the effect $j$. We experimentally verify this property and its impact by computing the Residual Concept Causal Effect i.e., the ratio between the Concept Causal Effect (CaCE) (Goyal et al., 2020) after and before blocking. The optimal value of this metric is zero, corresponding to perfect causal independence between $i$ and $j$ (i.e., an ideal de-biasing operation). The results show that, in Causal CGMs, blocking a variable in the causal graph always yields a perfect Residual Concept Causal Effect of zero across all datasets. In contrast, applying the same procedure in CBMs only reduces the average causal effect slightly, by just $3$ percentage points. CEMs not only fail to reduce the causal effect to zero but, in some cases, even increase the causal influence. These results underscore how Causal CGMs transparency enable users to manipulate the model’s decision-making process to achieve desired outcomes, as opposed to existing CBMs.
@@
checkmark dsprites celeba celeba
Black box N/A N/A N/A N/A
CBM $97.99{\scriptstyle\pm 5.64}$ $100.00{\scriptstyle\pm 0.70}$ $97.84{\scriptstyle\pm 2.13}$ $100.00{\scriptstyle\pm 0.00}$
CEM $102.58{\scriptstyle\pm 12.95}$ $100.00{\scriptstyle\pm 4.62}$ $106.00{\scriptstyle\pm 0.50}$ $100.00{\scriptstyle\pm 0.00}$
Causal CGM+CD $0.00{\scriptstyle\pm 0.00}$ $0.00{\scriptstyle\pm 0.00}$ $0.00{\scriptstyle\pm 0.00}$ N/A
Causal CGM $0.00{\scriptstyle\pm 0.00}$ $0.00{\scriptstyle\pm 0.00}$ $0.00{\scriptstyle\pm 0.00}$ $0.00{\scriptstyle\pm 0.00}$
@@ 5 Discussion
Causal CGMs present substantial advantages compared to the state of the art. Compared to most causal feature-attribution methods (e.g., (Chattopadhyay et al., 2019)), Causal CGMs focus on high-level human interpretable concepts. Causal CGMs differ from existing CBMs in their approach to intervention and causal relationships. The causal structure of the inferences that Causal CGMs deliver does not subsume weak forms of causal dependencies as in existing CBMs, where concepts are all direct causes of the target and causally independent of each other. This way, Causal CGMs enable human-in-the-loop corrections to mispredicted intermediate reasoning steps, boosting not just downstream accuracy after corrections but also the accuracy of the explanation provided for a specific instance. Closest to Causal CGMs in concept usage are post-hoc causal concept-based explainability techniques, like DiConStruct (de Oliveira Moreira et al., 2024) and conceptual counterfactual explanations (Abid et al., 2021). These methods build surrogate causal models to emulate a target black box model’s predictive behaviour. However, as Rudin (2019) notes, convergence in predictive behaviour does not ensure structural similarity in decision-making processes, making surrogate models as explanatory proxies questionable. Causally interpretable by-design architectures, such as Causal CGMs, do not suffer from this issue.
@@
Our method’s limitations mainly stem from inherent constraints in CBMs and causal reasoning. The quality of learnt causal graphs depends on dataset quality and annotations, where missing or noisy labels can lead to suboptimal graphs. Similarly to generalised PGMs, Causal CGMs can be easily unfolded when the final graph is acyclic. Variables in cycles can still be inferred, but require special unfolding techniques which might be explored in future works. Finally, a Causal CGM’s learnt graph does not necessarily represent the causal mechanisms of the data-generating process, but rather that of the model’s inference. This makes our models suitable to be verified and controlled but not necessarily to understand their training data’s distribution.
6 Conclusion
@@
Causal opacity represents a key open challenge at the intersection of deep learning, interpretability, and causality. Causal CGMs address this challenge by employing an architecture which makes the decision-making process causally transparent by design. This makes Causal CGMs reliable and verifiable compared to both usual DL architectures and standard (non-causal) CBMs. The results of our experiments show that Causal CGMs support the analysis of interventional and counterfactual scenarios—thereby improving the model’s causal interpretability and supporting the effective verification of its reliability and fairness—and enable human-in-the-loop corrections to mispredicted intermediate reasoning steps, boosting not just downstream accuracy after corrections, but also accuracy of the explanation provided for a specific instance. As a result, advancing this research line could significantly improve the reliability and verifiability of concept-based deep learning models, thus supporting their deployment in real-world applications.
Appendix A Causal explainability, causal opacity, and causal discovery
Deep Learning (DL) models have a pervasive impact on many areas of contemporary research and society (Baldi, 2021). Despite this success, there is growing concern about the widespread real-world application of DL, particularly in sensitive domains (Dignum, 2019). This is partly due to lack of causal explainability of these models, which undermine their robustness, fairness and generalisability in out-of-distribution context (Schölkopf et al., 2021). Causal explainability, in particular, is a multi-faced issue, which includes a variety of different, albeit related, problems (Termine & Primiero, 2024). One such problem is that of causal discovery and concerns the possibility of using a model to detect and understand the causal mechanisms of the data generating process (Pearl & Mackenzie, 2018; Pearl, 2019; Schölkopf et al., 2021; Kaddour et al., 2022). An equally important but distinct problem is that of causal opacity, which denotes the difficulty of users to grasp and understand the “hidden” causal structure that underlines the predictions delivered by a given model (see Figure 1(a)). This problem can be better understood in the light of Pearl’s framework of causality (Pearl, 2009; 2019), which relates causal understanding to the ability of an agent to answer what-if type of questions. In particular, Pearl identifies three different kinds of what-if questions organised in a hierarchy of levels. At the bottom level we have observational questions, which concern the actual state of affairs one can observe in data. At the intermediate level we find interventional questions, which concern the effects of intervening on the actual state of affairs and fix the value of some variable to a pre-defined one. Finally, the top level encompasses counterfactual questions, which concern an hypothetical state of affairs that could have been occured but it is not the actual one. The three levels provide a measure of the causal understanding that an agent has of a given target-system. Agents that answer only observational questions do not possess any causal understanding of the target system; agents with an intermediate level of causal understanding can answer interventional questions, whereas a complete causal understanding is required for counterfactual questions.
In this paper, we propose to involve Pearl’s hierarchy to define and measure the causal opacity/transparency of a DNN model based on the ability of its users to answer interventional and counterfactual questions concerning the structure of its inferential behaviour (e.g., “what happens if I fix the feature age to a value greater than $50$?” or “what the model’s prediction would have been if the feature age had taken a value greater than $50$ instead of lower then or equals to $50$?”). In this regard, complete causal transparency require that users can answer both interventional and counterfactual questions concerning the model’s inferential behaviour, whereas if only observational questions can be addressed (i.e., questions regarding the input-output associations of the model), then the model results completely causally opaque: this is notably the case of the majority of DNNs considered in contemporary AI research.
Notice that causal opacity is fundamentally distinct from the problem of causal discovery, which concerns the causal structure of the world and not that of a model’s inference. To clarify this point, consider a simple rules-based model as the one described in Fig. 7.
Figure 7: Rules-based Model $M_{1}$
This model is causally transparent as users can easily compute interventional and counterfactual questions concerning its inferential behaviour (e.g., “what if the age of the patient would had been $\geq 50$ instead of $<50$”?). However, causal transparency does not imply that the causal structure of the model’s inference (Fig.8(a)) resemble the causal structure of the world (Fig. 8(b)), which remains a very distinct animal.
AgeIncomeRisk of Hearth Attack
AgeIncomeRisk of Hearth Attack
Figure 8: A graphical representation of the causal structure of $M_{1}$’s inferential behaviour (a) as opposed to the real causal structure of the world (b). As $M_{1}$ uses the variable Income as a predictor of the variable Risk of Hearth Attack, a causal edge is present between the two variables in a. This edge is not present in b as no direct causal connection exists between the two variables in the real world.
To ensure that a model is causally transparent and matches the causal structure of the world, indeed, a suitable combination of causal transparency and causal discovery techniques is usually required. However, causal transparency per-sé is a fundamental requirement and sufficient for a variety of tasks, such as establishing the model’s fairness (especially within the popular framework of counterfactual fairness  (Kusner et al., 2017)).
AgeIncomeRisk of Hearth AttackGender
Figure 9: (a) Slightly modified version of $M_{1}$, and (b) a graphical representation of the causal structure of its inferential behaviour.
Consider the model in Fig. 9, which is a slightly modified version of the model depicted in Fig. 7. The causal transparency of the model makes it easy for users to verify the un-fairness of its inferential behaviour (stemming from the fact that the model uses the variable gender to predict the variable income), and implement suitable corrections (e.g. remove the causal link between gender and income).
Appendix B Derivation of Causal CGMs
In the absence of causal priors, Causal CGMs need to learn causal dependencies between high-level features that lead to the model becoming more effective in solving its designated task. As a result, we can formally describe Causal CGMs using a generalised probabilistic graphical model (PGM) that extends a traditional PGM by allowing cycles:
Definition B.1 (Generalized Causal Concept Graph Model).
Given an observed input feature $x$, a set of $k\in\mathbb{N}$ latent factors $u_{i}\in\mathcal{U}$ each associated with a high level interpretable variable $v_{i}\in\mathcal{V}$, a Generalized Causal Concept Graph Model is the generalised probabilistic graphical model (PGM) $G=(\mathcal{N},\mathcal{E})$ with nodes $\mathcal{N}=\{x\}\cup\mathcal{U}\cup\mathcal{V}$ and edges =$\mathcal{E}=\big{\{}(x,u_{i})\mid u_{i}\in\mathcal{U}\big{\}}\cup\big{\{}(u_{i},v_{i})\mid i\in\{1,\cdots,k\}\big{\}}\cup\{(v_{i},v_{j})\mid v_{i},v_{j}\in\mathcal{V},v_{i}\neq v_{j}\}$ which represents the joint conditional distribution $p(v,u\mid x)$:
[…]
The cyclical nature of Causal CGMs comes from the necessity to model all possible dependencies among variables $v_{i}$. Notice that the model is uniquely identified by the set of all conditional probability distributions corresponding to the arrows in the graph. Unfortunately, in generalised PGMs, the model does not easily factorise in terms of such distributions due to the cycles. To deal with cycles while maintaining the independencies induced by the graph structure, we can use an unfolding semantics for cyclical PGMs (Baier et al., 2022). This semantics is based on the choice of a “cutset” i.e., a specific set of nodes $\mathcal{Q}\subseteq\mathcal{N}$ in the PGM such that every cycle in the PGM contains at least one node in $\mathcal{Q}$. Intuitively, by unfolding the nodes in the cutset, all cycles are broken leaving us with a standard acyclical PGM. The consistency between the semantics of the original cyclical PGM and the unfolded acyclical PGM is only valid in the limit of infinite unfolding (Baier et al., 2022). However, when computing the likelihood of an observed complete set of variables $v_{i}\in\mathcal{V}$, modelling one single unfolding (i.e., a single transition) suffices for learning the conditional probability distributions among the variables in $\mathcal{V}$, as all the variables become conditionally independent on each other. As a result, we can define a Dissected Causal CGM as the one-step unfolding of the Causal CGM in Definition B.1:
Definition B.2 (Dissected Causal CGM).
Given a Causal CGM, let $\mathcal{V}^{\prime}=\mathcal{V}$ be the cutset. Then, the dissected Causal CGM $\mathcal{G}=(\mathcal{N}\cup\mathcal{V}^{\prime},\mathcal{E}_{\mathcal{V}^{\prime}})$ is an acyclic PGM obtained by extending the generalised PGM by (i) adding a copy of all cutset nodes $\mathcal{V}^{\prime}=\{v_{i}\mid v_{i}\in\mathcal{V}\}$, (ii) adding a new set of edges directed from parents of cutset’s nodes to the generated copies $\mathcal{V}^{\prime}$ i.e., $\mathcal{E}_{\mathcal{V}^{\prime}}=\{(a,b)\mid(a,b)\in\mathcal{E},b\in\mathcal{V}^{\prime}\}\cup\{(a,b)\mid(a,b)\in\mathcal{E},b\notin\mathcal{V}^{\prime}\}$, and (iii) defining an initial probability distribution for the new copies $p(v^{\prime}|u)$ given the latent variables. The resulting PGM, factorised as $p(v,v^{\prime},u\mid x)=p(v\mid v^{\prime},u)p(v^{\prime}\mid u)p(u\mid x)$, is:
[…]
Appendix C Proof of theorem 3.2
Proof.
Let us consider the graph $\mathcal{G}^{*}=(\mathcal{V}\cup{roots(\mathcal{G}^{\prime})},\{(v_{i},v_{j})\mid v_{i},v_{j}\in\mathcal{V},(v_{i},v_{j})\in\mathcal{E}_{\mathcal{G}^{\prime}}\}\cup{(v_{i}^{\prime},v_{j})|v_{i}^{\prime}\in roots(\mathcal{G}^{\prime})})$. We will first prove that this graph is isomorphic to $\mathcal{G}^{\prime}$ for all nodes $\mathcal{V}$ and then that the conditional probabilities are equivalent.
The graph $\mathcal{G}^{\prime}$ is defined with nodes $\mathcal{V}\cup\mathcal{V}^{\prime}$, where $\mathcal{V}$ represents the endogenous variables and $\mathcal{V}^{\prime}$ represents copies of the same high-level variables. Hence, the set of nodes $\mathcal{V}$ is the same for both graphs by definition. In $\mathcal{G}^{\prime}$, the edge set $\mathcal{E}_{\mathcal{G}^{\prime}}$ consists of edges $(v_{i}^{\prime},v_{j})$ between nodes in $\mathcal{V}^{\prime}$ and $\mathcal{V}$. Specifically, each $v_{i}^{\prime}\in\mathcal{V}^{\prime}$ is connected to $v_{j}\in\mathcal{V}$, where $v_{i}^{\prime}\neq v_{j}$. In $\mathcal{G}^{*}$, the edge set $\mathcal{E}_{\mathcal{G}^{*}}$ consists of:
  • edges $(v_{i},v_{j})$ between nodes in $\mathcal{V}$, which are present if $(v_{i},v_{j})\in\mathcal{E}_{\mathcal{G}^{\prime}}$,
  • edges $(v_{i}^{\prime},v_{j})$ where $v_{i}^{\prime}\in\text{roots}(G^{\prime})$, representing connections between the root nodes and their corresponding child nodes.
which represent a one-to-one correspondence between the edges in $\mathcal{E}_{\mathcal{G}^{\prime}}$ and $\mathcal{E}_{\mathcal{G}^{*}}$. Thus, the two graphs are isomorphic.
Now, we prove that for all nodes $v_{i}\in\mathcal{V}\setminus(\text{roots}(\mathcal{G}^{\prime})\cup\text{ch}(\text{roots}(\mathcal{G}^{\prime})))$, the conditional probabilities are equivalent: $p(v_{i}\mid\text{pa}_{\mathcal{V}^{\prime}}(v_{i}),u_{i};\theta_{f})=p(v_{i}\mid\text{pa}_{\mathcal{V}}(v_{i}),u_{i};\theta_{f})$. In $\mathcal{G}^{\prime}$, the conditional distribution of a node $v_{i}\in\mathcal{V}$ depends on its parents $\text{pa}_{\mathcal{V}^{\prime}}(v_{i})$, which are nodes in $\mathcal{V}^{\prime}$, as well as the latent factors $u_{i}$. Since $\mathcal{V}=\mathcal{V}^{\prime}$ by Def. 3.1 and the graphs $\mathcal{G}^{\prime}$ and $\mathcal{G}^{*}$ are isomorphic for all $\mathcal{V}$ nodes, the parent set $\text{pa}_{\mathcal{V}^{\prime}}(v_{i})$ in $\mathcal{V}^{\prime}$ corresponds directly to the parent set $\text{pa}_{\mathcal{V}}(v_{i})$ in $\mathcal{V}$, up to the bijection between the two sets. Under full observability (the training conditions of CGMs), conditional queries of the form $p(v|\text{ all but }v)$ are just inspections of the conditional probability table. As a result, $p(v_{i}\mid\text{pa}_{\mathcal{V}^{\prime}}(v_{i}),u_{i};\theta_{f})$ and $p(v_{i}\mid\text{pa}_{\mathcal{V}}(v_{i}),u_{i};\theta_{f})$ correspond to repeating the same query on the same conditional probability table. Therefore, for all nodes $v_{i}\in\mathcal{V}\setminus(\text{roots}(\mathcal{G}^{\prime})\cup\text{ch}(\text{roots}(\mathcal{G}^{\prime})))$, we can write: $p(v_{i}\mid\text{pa}_{\mathcal{V}^{\prime}}(v_{i}),u_{i};\theta_{f})=p(v_{i}\mid\text{pa}_{\mathcal{V}}(v_{i}),u_{i};\theta_{f})$. ∎
Appendix D Neural parametrization of Concept Graph Models
We can interpret the factors of a dissected Causal CGM as follows: $p(u_{j}\mid\mathbf{x};\theta_{\zeta})$ is the exogenous encoder, i.e., a deterministic distribution that is parametrised by a neural network $\zeta:X\to U$. In CBMs, this function represents the input encoder. The exogenous encoder $\zeta$ generates exogenous variables $u_{i}\in\mathcal{U}$ mapping raw input features $\mathbf{x}$ (e.g., an image’s pixels) to latent embeddings $\hat{\mathbf{u}}_{i}\in\mathbb{R}^{q},\ q\in\mathbb{N}$. In practice, this process mirrors the generation of context vectors in Concept Embedding Models to avoid information bottlenecks which may negatively affect the model’s accuracy (Espinosa Zarlenga et al., 2022). First, the encoder $\psi:X\rightarrow H$ maps raw features to a latent code $\mathbf{h}\in H$. Then, a pair of neural networks $\{\phi_{i}^{+},\phi_{i}^{-}\}$ map the latent code into two different embeddings whose concatenation $[\phi_{i}^{+}(\mathbf{h}),\phi_{i}^{-}(\mathbf{h})]^{T}$ corresponds to the exogenous variable $\mathcal{U}_{i}$ of the $i$-th concept:
In contrast, $p(v_{j}^{\prime}\mid u_{j};\theta_{s})$ is the copies predictor. This is the product of $k$ independent Bernoulli distributions whose logits are parameterised by a neural network $s:U\to V$. In CBMs, the composition of the exogenous encoder and the concept predictor is often called concept encoder $g=\zeta\circ s$. In causal methods, this function represents a (supervised) causal feature learner (Kaddour et al., 2022). The copies predictor $s$ generates endogenous copies $v_{i}^{\prime}\in\mathcal{V}$ from latent embeddings $\hat{\mathbf{u}}_{i}$. This is obtained using a neural network classifier $s:U\to V$ as a scoring function as in (Espinosa Zarlenga et al., 2022):
Finally, $p(v_{i}\mid\text{pa}_{\mathcal{V}^{\prime}}(v_{i}),u_{i};\theta_{f})$ is the endogenous predictor. This distribution is the product of $k$ independent Bernoulli distributions whose logits are parameterised by a neural network $f:V^{k}\times U\to V$. The input to this function $\text{pa}_{\mathcal{V}^{\prime}}(v_{i})$ (representing direct causal dependencies) is weighted by a learnable adjacency matrix $M\subseteq\mathbb{R}^{k\times k}$, where each learnable weight $a_{ij}$ models the strength of the dependency of $v_{i}$ from its parents $v_{j}^{\prime}$. In CBMs this function is called a task predictor (Koh et al., 2020) and intuitively represents the analogous of the structural equations that model causal mechanisms in SCMs (Kaddour et al., 2022). The endogenous predictor generates the endogenous variables $\hat{v}_{i}\in\mathcal{V}$ by considering exogenous variables $\hat{\mathbf{u}}_{j}$ and copies $\hat{v}_{j}^{\prime}$ with $j\neq i$. First, the function $\omega:V\times U\to\mathbb{R}^{z}$ generates endogenous embeddings $\hat{\mathbf{v}}_{j}^{\prime}$ using exogenous variables $\hat{\mathbf{u}}_{j}$ and copies $\hat{v}_{j}^{\prime}$, following (Espinosa Zarlenga et al., 2022). Then, all endogenous embeddings are weighted by the strength of the dependency $a_{ij}$ and aggregated using a deepset-like neural network $f_{i}:\mathbb{R}^{z\times k}\to[0,1]$ which maps endogenous embeddings to endogenous predictions:
(endogenous variables) $\quad\hat{v}_{i}=f_{i}\left(\{a_{ij}\hat{\mathbf{v}}_{j}^{\prime}\}_{j\in\{1,\dots,k\}}\right).$ (14)
In order to learn explicit structural equations, existing logic-based aggregation methods can be used from the concept literature (Barbiero et al., 2022; 2023). App. F describes in more detail their adaptation in Causal CGM.
Appendix E Unfolding CGMs
Notice how learning a DAG together with Definition B.2 allows to unfold a Causal CGM’s endogenous predictor applying a directed message passing on the associated structural causal model $\mathcal{M}$, ensuring that the values of endogenous variables are derived solely from the nodes that are their ancestors on the causal graph (see Figure 10). As a first step, we compute the exogenous variables for all nodes. We then predict the values of endogenous variables in root nodes in the learned DAG from their corresponding exogenous variables. Following this, we can generate the endogenous embeddings for root nodes and aggregate endogenous embeddings to compute the value of endogenous variables of each child node. We repeat this process until all leaf nodes of the graph are reached. We can obtain this by replacing in Eq. 13 the endogenous copies $\hat{v}_{j}^{\prime}$ with the parents of the endogenous variable $v_{i}$:
Note that this causal unrolling guarantees two key properties (see Figure 3(a) and Fig. 10): (1) modifying the value of a cause (parent node) will impact the effect (child node) in our model, (2) conversely, intervening upon an effect does not alter the cause. This is because, in our model, information flows sequentially following the graph, mirroring the fundamental nature of causal effects. Consequently, this layer not only facilitates the computation of task predictions but also enables the exploration of causal relationships through do-interventions and counterfactual analysis. In App. H.6 we explain why $1$-step unfolding of the causal graph during training is equivalent to any k-step ($k>1$) and empirically show related results.
[…]
Figure 10: Unfolded Causal CGM’s endogenous predictor.
Appendix F Architecture
Initialisation of adjacency matrix